An airline tracks data on its flight arrivals. Over the past six months, 65 flights on one route arrived early, 273 arrived on time, 218 were late, and 44 were cancelled. What is the probability that a flight is early? On time? Late? Cancelled? Are these outcomes mutually exclusive? what is the probability that a flight is either early or on time?
1
Expert's answer
2010-07-07T06:11:00-0400

The probability that a flight is early: 65/(65+273+218+44)=65/600= 0.108.
The probability that a flight is On time:273/600 = 0.455.
The probability that a flight is Late: 218/600 = 0.363.
The probability that a flight is Cancelled: 44/600 = 0.073.
The probability that a flight is either early or on time: (65+273)/600 = 0.563.
These outcomes are mutually exclusive.